the width of the rectangle is 5 feet and the diagonal is 9 feet. what is the area of the rectangle?
Answer by RAY100(1637)

drawing a picture shows a rectangle (l*w) with diagonal
given diag = 9, and width=5
using pythagorous
a^2+b^2=c^2
l^2+5^2=9^2
l^2 +25 =81
subt 25 both sides
l^2 = 56
take sq rt both sides
l=7.483
area is L*W= 7.483 *5=37.417